OpenStudy (anonymous):

If m∠BAD = 65, what is the m∠DCB?

OpenStudy (anonymous):

help

OpenStudy (anonymous):

Since AB = BD, angle BAD = angle BDA = 65 degrees. Equal sides have equal angles opposite them In a triangle,the sum of all angles is 180 degrees. Therefore, angle DBA = 180 - 65 -65 = 50 In a triangle, the external angle is equal to the sum of opposite interior angles, Therefore, angle BDC = angle BAD + angle ABD = 65 + 50 = 115 In triangle BDC, BD = DC, therefore angle DBC = angle DCB Therefore, angle DBC + angle DCB +angle BDC = 180 2*angle DCB = 180-115 angle DCB = 65/2 = 32.5 degrees